OpenStudy (anonymous):

What is 23 expressed as a difference of two perfect squares?

OpenStudy (ash2326):

\[\huge {23= 12^2-11^2}\]

OpenStudy (anonymous):

11 isn't a perfect square. it has to be the sum of two PERFECT squares.

OpenStudy (ash2326):

oh 23=144-121

OpenStudy (ash2326):

144 is square of 12 121 is square of 11

OpenStudy (anonymous):

and they are perfect squares?

OpenStudy (anonymous):

the question is: "Every odd prime can be expressed as the difference of two perfect squares in one and only one way. Find the way for 23."

OpenStudy (ash2326):

A perfect square no. is the result of product of a number by itself . 4 is a perfect square cause 2*2=4 similarly 121= 11*11 144=12*12

OpenStudy (anonymous):

awesome thanks so much !

OpenStudy (ash2326):

Welcome:D